# Unmasking Trees for Tabular Data

### v4

@@ Abstract
Despite much work on advanced deep learning and generative modeling techniques for tabular data generation and imputation, traditional methods have continued to win on imputation benchmarks. We herein present UnmaskingTrees, a simple method for tabular imputation (and generation) employing gradient-boosted decision trees which are used to incrementally unmask individual features. This approach offers state-of-the-art performance on imputation, and on generation given training data with missingness; and it has competitive performance on vanilla generation. To solve the conditional generation subproblem, we propose a tabular probabilistic prediction method, BaltoBot, which fits a balanced tree of boosted tree classifiers. Unlike older methods, it requires no parametric assumption on the conditional distribution, accommodating features with multimodal distributions; unlike newer diffusion methods, it offers fast sampling, closed-form density estimation, and flexible handling of discrete variables. We finally consider our two approaches as meta-algorithms, demonstrating in-context learning-based generative modeling with TabPFN.
@@ 1 Introduction
Given a tabular dataset, it is frequently desirable to impute missing values within that dataset, and to generate new synthetic examples. On data generation, recent work (Jolicoeur-Martineau et al., 2024b) (ForestDiffusion) has shown state-of-the-art results on data generation using gradient-boosted trees (Chen and Guestrin, 2016) trained on diffusion or flow-matching objectives, outperforming deep learning-based approaches. However, this approach tended to struggle on tabular imputation tasks, outperformed by MissForest (Stekhoven and Bühlmann, 2012), an older multiple imputation approach based on random forests (Breiman, 2001).
@@
Our two methods are in fact meta-algorithms that, in combination, can create a generative model out of any probabilistic binary classifier. To demonstrate this flexibility, we swap out XGBoost (Chen and Guestrin, 2016) for TabPFN (Hollmann et al., 2022). TabPFN is a deep learning model pretrained to perform in-context learning for tabular classification. While it has state-of-the-art classification benchmark performance (McElfresh et al., 2024), it currently does not perform regression tasks, nor does it inherently perform generative modeling (Ma et al., 2024). Constructing a generative model out of TabPFN (Hollmann et al., 2022) was first proposed in TabPFGen (Ma et al., 2024), which approximates the posterior from TabPFN-provided likelihoods by iteratively applying stochastic gradient Langevin dynamics (Welling and Teh, 2011). But unlike the previous work, ours requires only a few TabPFN forward-passes for each sample rather than many iterative data updates.
We showcase UnmaskingTrees on two tabular case studies, and on the benchmark of 27 tabular datasets presented by Jolicoeur-Martineau et al. (2024b). Most notably on this benchmark, our approach offers state-of-the-art performance on imputation and on generation given training data with missingness; and it has competitive performance on vanilla generation. We also demonstrate that BaltoBot is on its own a promising method for probabilistic prediction, showing its advantages on synthetic case studies and on a heavy-tailed sales forecasting benchmark.
Finally, we provide code with an easy-to-use sklearn-style API at https://github.com/calvinmccarter/unmasking-trees. In addition to being useful for practitioners, we hope our work sparks study within the tabular ML community about whether diffusion or autoregression is better for tabular data. Previous autoregressive tabular modeling methods, TabMT (Gulati and Roysdon, 2024) and DP-TBART (Castellon et al., 2023), use Transformer (Vaswani, 2017) models, making them less applicable for the GPU-poor. Our simple, efficient implementations of UnmaskingTrees and BaltoBot contribute to investigating this question.
@@ 2.1 UnmaskingTrees for tabular joint distribution modeling
UnmaskingTrees combines the gradient-boosted trees of ForestDiffusion (Jolicoeur-Martineau et al., 2024b) with the training objective of generalized autoregressive language modeling (Yang, 2019), inheriting the benefits of both. Consider a dataset with $N$ examples and $D$ features. We learn the joint distribution over $D$-dimensional example ${\mathbf{x}}$ by maximizing the expected log-likelihood with respect to all possible permutations of the factorization order,
@@
where $\sigma$ is a permutation drawn uniformly from ${\mathcal{U}}(G_{D})$, the permutation group on $D$ features; ${\mathbf{x}}_{\sigma(<t)}$ denotes all features that precede the $t$-th feature in the permuted sequence of features. If we were to have marginalized over permutations, we would have obtained a masked language modeling procedure with a randomly-sampled masking rate $r\sim{\mathcal{U}}(0,1)$ (Kitouni et al., 2023, 2024); such a procedure was previously shown to have benefits in combination with tabular Transformer models (Gulati and Roysdon, 2024) (TabMT).
For each example, we generate new training samples by randomly sampling an order over the features, then incrementally masking the features in that random order. Given duplication factor $K$, we repeat this process $K$ times with $K$ different random permutations, leading to a training dataset with $KND$ samples. Given this, we train XGBoost (Chen and Guestrin, 2016) models to predict each unmasked sample given the more-masked example derived from it, one per feature. We model categorical features via softmax-based classification with cross-entropy loss; our approach for continuous features is described in Section 2.2.
For both generation and imputation, we generate features of each sample in random order. For imputation rather than generation tasks, we begin by filling in each sample with the observed values, and run inference on the remaining unobserved features. Implementing this is very simple: it requires about 70 lines of Python code for training, and about 20 lines for inference.
@@ 2.2 BaltoBot for tabular probabilistic prediction
A key problem when autoregressively generating continuous data is that a regression model will attempt to predict the mean of a conditional distribution, whereas we would like it to sample from the possibly-multimodal conditional distribution. The simplest solution is to quantize continuous features into bins, because classification over histograms is inherently multimodal; TabMT (Gulati and Roysdon, 2024) did this with 1d k-Means clustering (Lloyd, 1982). Yet this not only destroys information within bins due to rounding, it also destroys information about the proximity among the ordered bins. Thus, it forces us to choose between a small number of quantization bins, yielding low resolution; or to choose a large number of bins, risking catastrophic errors due to overfitting and/or clumping of generated samples due to poor calibration. This not only limits performance, but also necessitates hyperparameter tuning (Gulati and Roysdon, 2024).
@@
At training and inference time, each XGBoost model within the meta-tree only sees examples that fall into its corresponding region of the output space. Thus, for a meta-tree with height $H$ (and thus $2^{H}$ models), each example is only passed as input to $H$ different models. While lower-level classifiers receive less data and are poorer quality, the magnitude of such errors are smaller due to our hierarchical partitioning approach. Furthermore, our singleton-bin technique allows us to adaptively generate discrete and even mixed-type variables, if these discrete outcomes are high-frequency relative to the total size of the data and to the height of the meta-tree. (Up to $2^{H}$ discrete outcomes can be produced by BaltoBot.) Finally, eschewing diffusion modeling enables us to perform closed-form conditional density estimation.
Algorithms for UnmaskingTrees and BaltoBot are given in Appendix Section A.
@@ 2.4 In-context learning-based generation with BaltoBoTabPFN and UnmaskingTabPFN
Within our flexible frameworks for joint and conditional modeling, TabPFN (Hollmann et al., 2022) can be used as a base learner for probabilistic prediction and generative modeling. For UnmaskingTabPFN joint modeling, a difficulty arises from TabPFN’s inability to handle inputs ${\mathbf{X}}_{\textrm{train}}$ with missing values (NaNs). To address this, we developed NanTabPFN, a wrapper for TabPFN that supports missingness in both training and test features. Based on each test query ${\mathbf{x}}_{\textrm{test}}$, we select row indices ${\mathcal{R}}$ and column indices ${\mathcal{C}}$ so that $[{\mathbf{X}}_{\textrm{train}}]_{{\mathcal{R}},{\mathcal{C}}}$ has no NaNs, using the following key idea. Consider a particular train example ${\mathbf{x}}_{\textrm{train}}$ and test query ${\mathbf{x}}_{\textrm{test}}$, with visible (non-missing) features denoted by sets ${\mathcal{V}}({\mathbf{x}}_{\textrm{train}})$ and ${\mathcal{V}}({\mathbf{x}}_{\textrm{test}})$. We can maximize the number of utilized features, while also ensuring that TabPFN receives no NaNs, by restricting the set of columns to those observed for the test query, ${\mathcal{C}}:={\mathcal{V}}({\mathbf{x}}_{\textrm{test}})$, then choosing training examples :=${\mathcal{R}}:=\{i|{\mathcal{C}}\subseteq{\mathcal{V}}({\mathbf{x}}^{(i)}_{\textrm{train}})\}$. In practice, our procedure is more complicated, because the above choices may result in either empty ${\mathcal{C}}$ or empty ${\mathcal{R}}$. If ${\mathcal{R}}$ is empty, we incrementally set random features of ${\mathbf{x}}_{\textrm{test}}$ to missing until we are able to obtain a non-empty training set. If ${\mathcal{C}}$ is empty, we introduce a new all-1s feature to both ${\mathbf{X}}_{\textrm{train}}$ and ${\mathbf{x}}_{\textrm{test}}$.
@@ 3 Results
We evaluate UnmaskingTrees on two case studies (Section 3.1) and on a tabular benchmark of 27 datasets (Section 3.2). We then evaluate BaltoBot and BaltoBoTabPFN on tabular probabilistic prediction case studies (Section 3.3) and on a sales forecasting dataset (Section 3.4). Results were obtained always using the default hyperparameters: output tree height of 4, and duplication factor $K=50$. These hyperparameter values were tuned on the Two Moons and Iris case studies, then applied without further tuning to the remaining experiments, because hyperparameter tuning is no fun at all. XGBoost hyperparameters were set to their defaults. Experiments were performed on a iMac (21.5-inch, Late 2015) with 2.8GHz Intel Core i5 processor and 16GB memory.
Overall, UnmaskingTrees (using BaltoBot) has state-of-the-art performance on imputation and on generation after training on incomplete data; and it has competitive performance on vanilla tabular generation scenarios. We further demonstrate the benefits of BaltoBot and BaltoBoTabPFN when evaluated in their own right for probabilistic prediction.
@@ 3.2 Benchmarking UnmaskingTrees on 27 tabular datasets
Here, we add UnmaskingTrees to the benchmark of 8 imputation methods on 27 public datasets, evaluated according to 9 metrics, developed by Jolicoeur-Martineau et al. (2024b) for evaluating tabular imputation and generation methods. This benchmark primarily contains smaller-sized (with $640103\leq N\leq 20{,}$ 640 and $4\leq D\leq 90$) datasets, which our approach is especially geared towards. Namely, we compare our approach against Forest-VP Jolicoeur-Martineau et al. (2024b), as well as k-NN imputation (Troyanskaya et al., 2001), ICE (Buck, 1960), MICE-Forest (Van Buuren et al., 1999; Wilson et al., 2022), MissForest (Stekhoven and Bühlmann, 2012), Softimpute (Hastie et al., 2015), minibatch Sinkhorn optimal transport (Muzellec et al., 2020), and generative adversarial nets (GAIN) (Yoon et al., 2018).  [^1] We do not add TabMT (Gulati and Roysdon, 2024) and TabPFGen (Ma et al., 2024) to the benchmark because no code was provided. We do not add UnmaskingTabPFN because of out-of-memory errors on our machine. We follow Jolicoeur-Martineau et al. (2024b) in computing the per-dataset rank of each method relative to other methods, then reporting the average over 27 datasets. For all methods other than our own, we compute ranks by reusing the raw scores provided in Jolicoeur-Martineau et al. (2024b)’s code repository.
@@
Table 1: Tabular data imputation (27 datasets, 3 experiments per dataset, 10 imputations per experiment) with 20% missing. Shown are averaged rank over all datasets and experiments (standard-error). Overall best is highlighted; better of Forest-VP versus ours is boldface blue. Metrics are Minimum and Average mean-absolute error (MinMAE and AvgMAE) to ground-truth, Wasserstein distance to train and test dataset distributions ($W_{train}$ and $W_{test}$), Mean Absolute Deviation (MAD) around the median/mode (for diversity), $R^{2}$ and $F_{1}$ for downstream regression / classification problems, and percent bias $P_{bias}$ and confidence interval coverage rate $Cov_{rate}$ for statistical inferences.
@@
We next repeat the experimental setup of Jolicoeur-Martineau et al. (2024b) for evaluating tabular generation methods. For tabular generation, using the same 27 datasets, Jolicoeur-Martineau et al. (2024b) benchmark their methods (Forest-VP and Forest-Flow) against 6 other methods, namely, Gaussian Copula (Joe, 2014), tabular variational autoencoding (TVAE) (Xu et al., 2019), two conditional generative adversarial net methods (CTGAN (Xu et al., 2019) and CTAB-GAN+ (Zhao et al., 2021)), and two other tabular diffusion methods (STaSy (Kim et al., 2022) and TabDDPM (Kotelnikov et al., 2023)). These are evaluated with 9 metrics, in the vanilla fully-observed setting and in the synthetically-induced 20% missing completely at random (MCAR) setting.

### v5

# algorithmic

@@ Abstract
Despite much work on advanced deep learning and generative modeling techniques for tabular data generation and imputation, traditional methods have continued to win on imputation benchmarks. We herein present UnmaskingTrees, a simple method for tabular imputation (and generation) employing gradient-boosted decision trees which are used to incrementally unmask individual features. On a benchmark for “out-of-the-box” performance on 27 small tabular datasets, UnmaskingTrees offers leading performance on imputation; state-of-the-art performance on generation given data with missingness; and competitive performance on vanilla generation given data without missingness. To solve the conditional generation subproblem, we propose a tabular probabilistic prediction method, BaltoBot, which fits a balanced tree of boosted tree classifiers. Unlike older methods, it requires no parametric assumption on the conditional distribution, accommodating features with multimodal distributions; unlike newer diffusion methods, it offers fast sampling, closed-form density estimation, and flexible handling of discrete variables. We finally consider our two approaches as meta-algorithms, demonstrating in-context learning-based generative modeling with TabPFN.
@@ 1 Introduction
Given a tabular dataset, it is frequently desirable to impute any missing values within that dataset, and to generate new synthetic examples. Due to the prevalence of missingness in tabular datsets, imputation has been a long-standing task in statistics and machine learning (Little and Rubin, 2019). In particular, multiple imputation methods, which produce multiple samples from the estimated conditional distribution of missing features, have proved advantageous for downstream inferential and prediction tasks (Rubin, 1996). Multiple imputation has also been used as a subroutine for counterfactual estimation (Kreindler and Lumsden, 2016; Yoon et al., 2018b) and domain adaptation (Ragab et al., 2023; McCarter, 2024). Synthetic data generation for tabular data has also seen recent interest, with applications in addressing data imbalance (Van Breugel et al., 2021; Kim et al., 2022b) and in preserving privacy (Kotelnikov et al., 2023; Gulati and Roysdon, 2024).
Imputation and generation are closely related tasks. Multiple imputation can be seen as a form of conditional generation, where the partitioning between output variables and input variables is not known in advance. Generation is then a special case of imputation where the set of observed conditioning variables is empty. Furthermore, due to the aforementioned prevalance of missingness in tabular data, generation methods also frequently need to be able to handle missingness at training time.
In this work, we are primarily focused on generation and imputation methods for users with limited data and computing resources. On data generation, recent work (Jolicoeur-Martineau et al., 2024b) (ForestDiffusion) has shown leading results on data generation using gradient-boosted trees (Chen and Guestrin, 2016) trained on diffusion or flow-matching objectives, outperforming deep learning-based approaches, particularly on smaller datasets. However, this approach tended to struggle on tabular imputation tasks, outperformed by MissForest (Stekhoven and Bühlmann, 2012), an older multiple imputation approach based on random forests (Breiman, 2001).
@@
Our two methods are in fact meta-algorithms that, in combination, can create a generative model out of any probabilistic binary classifier. To demonstrate this flexibility, we swap out XGBoost (Chen and Guestrin, 2016) for TabPFN (Hollmann et al., 2022). TabPFN is a deep learning model pretrained to perform in-context learning for tabular classification. While it has state-of-the-art classification benchmark performance (McElfresh et al., 2024), it does not perform regression tasks, nor does it inherently perform generative modeling. Constructing a generative model out of TabPFN (Hollmann et al., 2022) was first proposed in TabPFGen (Ma et al., 2024), which approximates the posterior from TabPFN-provided likelihoods by iteratively applying stochastic gradient Langevin dynamics (Welling and Teh, 2011). But unlike the previous work, ours requires only a few TabPFN forward-passes for each sample rather than many iterative data updates.
We showcase UnmaskingTrees on two tabular case studies, and on the benchmark of 27 tabular datasets presented by Jolicoeur-Martineau et al. (2024b). On this benchmark for “out-of-the-box” performance on small tabular datasets, our approach offers leading performance on imputation and state-of-the-art performance on generation given data with missingness; and it has competitive performance on vanilla generation without missingness. We also demonstrate that BaltoBot is on its own a promising method for probabilistic prediction, showing its advantages on synthetic case studies and on a heavy-tailed sales forecasting benchmark.
Finally, we provide code with an easy-to-use sklearn-style API at https://github.com/calvinmccarter/unmasking-trees. In addition to being useful for practitioners, we hope our work sparks study within the tabular ML community about whether diffusion or autoregression is better for tabular data. Previous autoregressive tabular modeling methods, TabMT (Gulati and Roysdon, 2024) and DP-TBART (Castellon et al., 2023), use Transformer (Vaswani, 2017) models, making them less applicable for the GPU-poor; they also lack publicly-available implementations. Our simple, efficient implementations of UnmaskingTrees and BaltoBot contribute to investigating this question.
@@ 2.1 UnmaskingTrees for tabular joint distribution modeling
UnmaskingTrees combines gradient-boosted trees with the training objective of generalized autoregressive language modeling (Yang, 2019), inheriting the benefits of both. Consider a dataset with $N$ samples and $D$ features. We learn the joint distribution over $D$-dimensional sample ${\mathbf{x}}$ by maximizing the expected log-likelihood with respect to all possible permutations of the factorization order,
@@
where $\sigma$ is a permutation drawn uniformly from ${\mathcal{U}}(G_{D})$, the permutation group on $D$ features; ${\mathbf{x}}_{\sigma(<t)}$ denotes all features that precede the $t$-th feature in the permuted sequence of features. If we were to have marginalized over permutations, we would have obtained a masked language modeling procedure with a randomly-sampled masking rate $r\sim{\mathcal{U}}(0,1)$ (Liao et al., 2020; Kitouni et al., 2023; 2024); such a procedure was previously shown to have benefits in combination with tabular Transformer models (Gulati and Roysdon, 2024) (TabMT).
For each sample, we generate new training samples by randomly sampling an order over the features, then incrementally masking the features in that random order. Given duplication factor $K$, we repeat this process $K$ times with $K$ different random permutations, leading to a training dataset with $KND$ samples. Given this, we train XGBoost (Chen and Guestrin, 2016) models to predict each unmasked sample given the more-masked sample derived from it, one per feature. We model categorical features via softmax-based classification with cross-entropy loss; our approach for non-categorical features is described in Section 2.2.
For both generation and imputation, we generate features of each sample in random order. For imputation rather than generation tasks, we begin by filling in each sample with the observed values, and run inference on the remaining unobserved features. Implementing this is very simple: it requires about 70 lines of Python code for training, and about 20 lines for inference. The training algorithm for UnmaskingTrees is given in Algorithm 1.
Algorithm 1 Unmasking Trees training
0:  dataset ${\mathbf{X}}\in{\mathbb{R}}^{N\times D}$; duplication factor $K$.
1:  {# Build self-supervised training set}
3:  for $k=1,\dots,K$ do
4:     for $n=1,\dots,N$ do
5:        Draw random permutation $\sigma$ from ${\mathcal{U}}(G_{D})$
6:        Set $\bm{x}:={\mathbf{X}}_{n,:}$ and $\bm{y}:={\mathbf{X}}_{n,:}$.
7:        for $d=1,\dots,D$ do
[…]
10:        end for
11:     end for
12:  end for
13:  {# Train conditional generation models}
14:  for $d=1,\dots,D$ do
15:     if feature $d$ in ${\mathbf{X}}$ is a categorical feature then
16:        Train XGBClassifier on $([{\mathbf{X}}_{\textrm{train}}]_{:,j\neq d},[{\mathbf{Y}}_{\textrm{train}}]_{:,d})$.
17:     else
18:        Run BaltoBot with $([{\mathbf{X}}_{\textrm{train}}]_{:,j\neq d},[{\mathbf{Y}}_{\textrm{train}}]_{:,d})$.
19:     end if
20:  end for
@@ 2.2 BaltoBot for tabular probabilistic prediction
In diffusion models, predicting the score function can be framed as a regression problem where the model learns to estimate the conditional mean. However, a key problem when switching to autoregressively generating continuous data is that this regression approach will attempt to predict the mean of a conditional distribution, whereas we would like the model to sample from the possibly-multimodal conditional distribution. The simplest solution is to quantize continuous features into bins, because classification over histograms is inherently multimodal; TabMT (Gulati and Roysdon, 2024) did this with 1d k-Means clustering (Lloyd, 1982). Yet this not only destroys information within bins due to rounding, it also destroys information about the proximity among the ordered bins. Thus, it forces us to choose between a small number of quantization bins, yielding low resolution; or to choose a large number of bins, risking catastrophic errors due to overfitting and/or clumping of generated samples due to poor calibration. This not only limits performance, but also necessitates hyperparameter tuning (Gulati and Roysdon, 2024).
@@
Conceptually, our proposed approach has four advantages. First, at training and inference time, each XGBoost model within the meta-tree only sees samples that fall into its corresponding region of the output space. Thus, for a meta-tree with height $H$ (and thus $2^{H}$ models), each sample is only passed as input to $H$ different models. While lower-level classifiers receive less data and are poorer quality, the magnitude of such errors are smaller due to our hierarchical partitioning approach. Second, our singleton-bin technique allows us to adaptively generate discrete and mixed-type variables, if the discrete outcome is frequent relative to the total number of training samples and to the size of the meta-tree. (Up to $2^{H}$ discrete outcomes can be produced by BaltoBot.) Third, our adoption of KDI instead of KMeans for feature quantization is beneficial because tabular data often has features which are irregular or highly skewed. KMeans clustering is based on mixture modeling with equal-sized Gaussian components; in constrast, KDI is a 1d density-based clustering method looks for local minima after applying a smoothing transformation. KDI was previously shown (McCarter, 2023) to be better at discretizing such irregular features than KMeans, so we propose adopting it here for generative modeling. Finally, eschewing diffusion modeling enables us to perform closed-form conditional density estimation.
The training and inference algorithms for BaltoBot are given in Algorithms 2 and 3, respectively.
Algorithm 2 BaltoBot training
1:  if H = 0 or unique$({\mathbf{y}})=C$ for some constant $C$ then
2:     Save $\texttt{bounds}:=(\min({\mathbf{y}}),\max({\mathbf{y}}))$.
3:  else
4:     Obtain split point $p$ from KDI quantization on ${\mathbf{y}}$.
8:  end if
Algorithm 3 BaltoBot inference
0:  input query ${\mathbf{x}}\in{\mathbb{R}}^{D}$; trained BaltoBot model.
1:  if bounds is defined then
2:     Sample uniformly from $U(\texttt{bounds})$.
3:     Return.
4:  else
5:     Obtain prediction from XGBoost binary classifier.
6:     if prediction = left-child then
7:        Run inference on “left-child” BaltoBot with input query ${\mathbf{x}}$.
8:     else if prediction = right-child then
10:     end if
11:  end if
@@ 2.3 Computational complexity
We observe that MissForest and MICE-Forest require training $DT$ models, where number of iterations $T$ is determined by a stopping criterion that tests for convergence. Assuming that dataset size does not affect the number of required iterations, these methods each have complexity $O(ND^{2})$.
@@ 2.4 In-context learning-based generation with BaltoBoTabPFN and UnmaskingTabPFN
Within our flexible frameworks for joint and conditional modeling, TabPFN (Hollmann et al., 2022) can be used as a base learner for probabilistic prediction and generative modeling. For UnmaskingTabPFN joint modeling, a difficulty arises from TabPFN’s inability to handle inputs ${\mathbf{X}}_{\textrm{train}}$ with missing values (NaNs). To address this, we developed NanTabPFN, a wrapper for TabPFN that supports missingness in both training and test features. Based on each test query ${\mathbf{x}}_{\textrm{test}}$, we select row indices ${\mathcal{R}}$ and column indices ${\mathcal{C}}$ so that $[{\mathbf{X}}_{\textrm{train}}]_{{\mathcal{R}},{\mathcal{C}}}$ has no NaNs, using the following key idea. Consider a particular train sample ${\mathbf{x}}_{\textrm{train}}$ and test query ${\mathbf{x}}_{\textrm{test}}$, with visible (non-missing) features denoted by sets ${\mathcal{V}}({\mathbf{x}}_{\textrm{train}})$ and ${\mathcal{V}}({\mathbf{x}}_{\textrm{test}})$. We can maximize the number of utilized features, while also ensuring that TabPFN receives no NaNs, by restricting the set of columns to those observed for the test query, ${\mathcal{C}}:={\mathcal{V}}({\mathbf{x}}_{\textrm{test}})$, then choosing training samples :=${\mathcal{R}}:=\{i|{\mathcal{C}}\subseteq{\mathcal{V}}({\mathbf{x}}^{(i)}_{\textrm{train}})\}$. In practice, our procedure is more complicated, because the above choices may result in either empty ${\mathcal{C}}$ or empty ${\mathcal{R}}$. If ${\mathcal{R}}$ is empty, we incrementally set random features of ${\mathbf{x}}_{\textrm{test}}$ to missing until we are able to obtain a non-empty training set. If ${\mathcal{C}}$ is empty, we introduce a new all-1s feature to both ${\mathbf{X}}_{\textrm{train}}$ and ${\mathbf{x}}_{\textrm{test}}$.
@@ 3 Results
We evaluate UnmaskingTrees on two case studies (Section 3.1) and on a tabular benchmark of 27 datasets (Section 3.2). We then evaluate BaltoBot and BaltoBoTabPFN on tabular probabilistic prediction case studies (Section 3.3) and on a sales forecasting dataset (Section 3.4). Results were obtained always using our method’s default hyperparameters: BaltoBot tree height of 4, and duplication factor $K=50$. These hyperparameter values were tuned on the Two Moons and Iris case studies, then applied without further tuning to the remaining experiments, because tuning is problematic for users with limited computing and/or data resources. (The tree height $H$ was tuned on Two Moons and Iris by increasing $H$ until we saw that the resulting plots stopped showing visible improvement.) XGBoost hyperparameters were set to their defaults. Experiments were performed on a iMac (21.5-inch, Late 2015) with 2.8GHz Intel Core i5 processor and 16GB memory.
Overall, UnmaskingTrees (using BaltoBot) has leading performance on imputation and state-of-the-art performance on generation after training on incomplete data; and it has competitive performance on vanilla tabular generation scenarios. We further demonstrate the benefits of BaltoBot and BaltoBoTabPFN when evaluated in their own right for probabilistic prediction.
@@ 3.2 Benchmarking UnmaskingTrees on 27 tabular datasets
Here, we add UnmaskingTrees to the benchmark of 8 imputation methods on 27 public datasets, evaluated according to 9 metrics, developed by Jolicoeur-Martineau et al. (2024b) for evaluating tabular imputation and generation methods. This benchmark primarily contains smaller-sized (with $640103\leq N\leq 20{,}$ 640 and $4\leq D\leq 90$) datasets, which our approach is especially geared towards. Namely, we compare our approach against Forest-VP Jolicoeur-Martineau et al. (2024b), as well as k-NN imputation (Troyanskaya et al., 2001), ICE (Buck, 1960), MICE-Forest (Van Buuren et al., 1999; Wilson et al., 2022), MissForest (Stekhoven and Bühlmann, 2012), Softimpute (Hastie et al., 2015), minibatch Sinkhorn optimal transport (Muzellec et al., 2020), and generative adversarial nets (GAIN) (Yoon et al., 2018a).  [^1] We do not add TabMT (Gulati and Roysdon, 2024) and TabPFGen (Ma et al., 2024) to the benchmark because no code was provided. We do not add UnmaskingTabPFN because of out-of-memory errors on our machine. We follow Jolicoeur-Martineau et al. (2024b) in computing the per-dataset rank of each method relative to other methods, then reporting the average over 27 datasets. For all methods other than our own, we compute ranks by reusing the raw scores provided in Jolicoeur-Martineau et al. (2024b)’s code repository.
@@
Table 1: Tabular data imputation (27 datasets, 3 experiments per dataset, 10 imputations per experiment) with 20% missing. Shown are averaged rank over all datasets and experiments (standard-error). Overall best is highlighted; better of Forest-VP versus ours is boldface blue.
@@
We next repeat the experimental setup of Jolicoeur-Martineau et al. (2024b) for evaluating tabular generation methods. For tabular generation, using the same 27 datasets, Jolicoeur-Martineau et al. (2024b) benchmark their methods (Forest-VP and Forest-Flow) against 6 other methods, namely, Gaussian Copula (Joe, 2014), tabular variational autoencoding (TVAE) (Xu et al., 2019), two conditional generative adversarial net methods (CTGAN (Xu et al., 2019) and CTAB-GAN+ (Zhao et al., 2021)), and two other tabular diffusion methods (STaSy (Kim et al., 2022a) and TabDDPM (Kotelnikov et al., 2023)). These are evaluated with 9 metrics, in the vanilla fully-observed setting and in the synthetically-induced 20% missing completely at random (MCAR) setting.